# 7B Fully Open Source Moxin-LLM – From Pretraining to GRPO-based Reinforcement Learning Enhancement

### v4

@@ Abstract
Recently, Large Language Models (LLMs) have undergone a significant transformation, marked by a rapid rise in both their popularity and capabilities. Leading this evolution are proprietary LLMs like GPT-4 and GPT-o1, which have captured widespread attention in the AI community due to their remarkable performance and versatility. Simultaneously, open-source LLMs, such as LLaMA and Mistral, have made great contributions to the ever-increasing popularity of LLMs due to the ease to customize and deploy the models across diverse applications. Although open-source LLMs present unprecedented opportunities for innovation and research, the commercialization of LLMs has raised concerns about transparency, reproducibility, and safety. Many open-source LLMs fail to meet fundamental transparency requirements by withholding essential components like training code and data, and some use restrictive licenses whilst claiming to be “open-source,” which may hinder further innovations on LLMs. To mitigate this issue, we introduce Moxin 7B, a fully open-source LLM developed in accordance with the Model Openness Framework (MOF), a ranked classification system that evaluates AI models based on model completeness and openness, adhering to principles of open science, open source, open data, and open access. Our model achieves the highest MOF classification level of “open science” through the comprehensive release of pre-training code and configurations, training and fine-tuning datasets, and intermediate and final checkpoints, aiming to make continuous commitments to fully open-source LLMs. The pre-training cost is about \$160,000. After pre-training and obtaining the base model, we finetune the Moxin Base model with SOTA post-training framework and instruction data to obtain Moxin Instruct model. To improve the reasoning capability, we further finetune our Instruct model with chain-of-thought data distilled from DeepSeek R1, and then use Group Relative Policy Optimization (GRPO), an efficient and effective reinforcement learning algorithm following DeepSeek R1, to finetune our model, leading to the Moxin Reasoning model. Experiments show that our models achieve superior performance in various evaluations such as zero-shot evaluation, few-shot evaluation, and CoT evaluation, compared with other base models and instruct models. The outstanding performance of our Reasoning model demonstrates the effectiveness of RL for small LLMs such as 7B models. Besides our open pretraining with released base model, data, code, etc., in our post-training including the instruction finetuning and CoT finetuning, we adopt open-source training frameworks with available data, code, and configurations. We release our instruct model and reasoning model, along with the available data and code to derive these models.
@@
Base model: https://huggingface.co/moxin-org/moxin-llm-7b
Instruct model: https://huggingface.co/moxin-org/moxin-Instruct-7b
Reasoning model: https://huggingface.co/moxin-org/moxin-Reasoning-7b
@@ 1 Introduction
We also finetune the Moxin Base model with SOTA post-training framework and instruction data to obtain Moxin Instruct model. To improve the reasoning capability, we further finetune our model with chain-of-thought data distilled from DeepSeek R1, and then use Group Relative Policy Optimization, an efficient and effective reinforcement learning algorithm following DeepSeek R1, to finetune our model, leading to the Moxin Reasoning model. Experiments show that our models achieve superior performance in various evaluations such as zero-shot evaluation, few-shot evaluation, and CoT evaluation, compared with other instruct and reasoning models. The outstanding performance of our Reasoning model demonstrates the effectiveness of RL for small LLMs such as 7B models.
Besides our open pretraining with released base model, data, code, etc., in our post-training including the instruction finetuning and CoT finetuning, we adopt open-source training frameworks with available data, code, and configurations. We release our instruct model and reasoning model, along with the available data and code to derive these models. Our homepage is https://github.com/moxin-org/Moxin-LLM. We summarize our contributions below.
@@ 6.4 CoT Evaluation
[h] Our developed models and their names in our releases. Developed Models Names in Releases Moxin-7B-Enhanced Moxin-7B-Base Moxin-7B-SFT Moxin-7B-DPO Moxin-7B-Instruct Moxin-7B-DPO-II Moxin-7B-RL-DeepScaleR Moxin-7B-Reasoning Moxin-7B-RL-AReal
@@ 6.5 Generation Examples
We demonstrate some text generation examples from our Moxin Chat 7B and Llama2 Chat 7B in Table 6.5 and 6.5. Moxin Chat 7B is developed by finetune the Moxin base model on the Tulu2 dataset. The response of Moxin Chat 7B is more accurate and concise.

### v5

# 7B Fully Open Source Moxin-LLM/VLM – From Pretraining to GRPO-based Reinforcement Learning Enhancement

@@ Abstract
Recently, Large Language Models (LLMs) have undergone a significant transformation, marked by a rapid rise in both their popularity and capabilities. Leading this evolution are proprietary LLMs like GPT-4 and GPT-o1, which have captured widespread attention in the AI community due to their remarkable performance and versatility. Simultaneously, open-source LLMs, such as LLaMA and Mistral, have made great contributions to the ever-increasing popularity of LLMs due to the ease to customize and deploy the models across diverse applications. Although open-source LLMs present unprecedented opportunities for innovation and research, the commercialization of LLMs has raised concerns about transparency, reproducibility, and safety. Many open-source LLMs fail to meet fundamental transparency requirements by withholding essential components like training code and data, and some use restrictive licenses whilst claiming to be “open-source,” which may hinder further innovations on LLMs. To mitigate this issue, we introduce Moxin 7B, a fully open-source LLM developed in accordance with the Model Openness Framework (MOF), a ranked classification system that evaluates AI models based on model completeness and openness, adhering to principles of open science, open source, open data, and open access. Our model achieves the highest MOF classification level of “open science” through the comprehensive release of pre-training code and configurations, training and fine-tuning datasets, and intermediate and final checkpoints, aiming to make continuous commitments to fully open-source LLMs. The pre-training cost is about \$160,000. After pre-training and obtaining the base model, we finetune the Moxin Base model with SOTA post-training framework and instruction data to obtain Moxin Instruct model. To improve the reasoning capability, we further finetune our Instruct model with chain-of-thought data distilled from DeepSeek R1, and then use Group Relative Policy Optimization (GRPO), an efficient and effective reinforcement learning algorithm following DeepSeek R1, to finetune our model, leading to the Moxin Reasoning model. Moreover, based on an open-source vision language model (VLM) framework, we develop and release our VLM with our Moxin model as the LLM backbone. Experiments show that our models achieve superior performance in various evaluations such as zero-shot evaluation, few-shot evaluation, and CoT evaluation, compared with other base models and instruct models. The outstanding performance of our Reasoning model demonstrates the effectiveness of RL for small LLMs such as 7B models. Our VLM outperforms other VLM models or LLM backbones. Besides our open pretraining with released base model, data, code, etc., in our post-training including the instruction finetuning, CoT finetuning, and VLM finetuning, we adopt open-source training frameworks with available data, code, and configurations. We release our instruct model, reasoning model, and VLM model, along with the available data and code to derive these models.
@@
Base model: https://huggingface.co/moxin-org/Moxin-7B-LLM
Instruct model: https://huggingface.co/moxin-org/Moxin-7B-Instruct
Reasoning model: https://huggingface.co/moxin-org/Moxin-7B-Reasoning
VLM model: https://huggingface.co/moxin-org/Moxin-7B-VLM
@@ 1 Introduction
We also finetune the Moxin Base model with SOTA post-training framework and instruction data to obtain Moxin Instruct model. To improve the reasoning capability, we further finetune our model with chain-of-thought data distilled from DeepSeek R1, and then use Group Relative Policy Optimization, an efficient and effective reinforcement learning algorithm following DeepSeek R1, to finetune our model, leading to the Moxin Reasoning model. Moreover, we develop and release our vision language model (VLM) based on our Moxin model as the LLM backbone with an open-source VLM framework. Experiments show that our models achieve superior performance in various evaluations such as zero-shot evaluation, few-shot evaluation, and CoT evaluation, compared with other instruct and reasoning models. The outstanding performance of our Reasoning model demonstrates the effectiveness of RL for small LLMs such as 7B models. Our VLM outperforms other VLM models or LLM backbones.
Besides our open pretraining with released base model, data, code, etc., in our post-training including the instruction finetuning, CoT finetuning, and VLM finetuning, we adopt open-source training frameworks with available data, code, and configurations. We release our instruct model, reasoning model, and VLM model, along with the available data and code to derive these models. Our homepage is https://github.com/moxin-org/Moxin-LLM. We summarize our contributions below.
@@
  • Moreover, we develop and release our VLM based on our Moxin model as the LLM backbone. Based on the open-source VLM framework Prismatic VLMs, we train our VLM on fully open-source datasets with Moxin as the LLM backbone and DINOv2 & SigLIP as the visual backbone. Our experiments demonstrate that our VLM outperforms other VLM models or LLM backbones.
@@ 5 Chain-of-Thought Enhancement
6 Vision Language Model based on Moxin
Vision language models (VLMs) [111, 112, 113, 114], which leverage LLMs, have achieved significant breakthroughs, enabling sophisticated vision-language dialogues and interactions. To develop VLM, we adopt the Prismatic VLMs framework [115] to train our VLM based on our Moxin model. Specifically, for the visual backbone with image processing and visual representations, we adopt DINOv2 [116] and SigLIP [117], and fuse their features to provide significant gains. For the LLM part, we adopt our Moxin-7B-Base as the LLM backbone.
Following the hypotheses in [118] and similar work, the DINOv2 features provide features that capture low-level spatial properties of an image, augmenting the higher-level “semantic” properties captured by vision-language contrastive models. Typically, the visual module such as (CLIP and SigLIP) trained with the vision-language contrastive objective leads to much better performance. Furthermore, SigLIP contains internet-sourced images from multiple sources (e.g., sketches, diagrams, animated graphics, etc.) which are not in ImageNet or in the DINOv2 pretraining data. DINOv2 and SigLIP are good complements for the image processing and visual representations.
For the model architecture, we adopt the general architecture used by many recent VLMs, such as LLaVa, Qwen-VL, and PaLI-3 [119, 111, 120]. These architectures use a (pretrained) visual backbone to map an input image to a sequence of patch features that are then projected individually into the embedding space of an LM. Specifically, a VLM takes as input an image and text prompt tokens with arbitrary sequence length. These inputs are then fed to the following components: 1) a visual representation backbone, 2) a vision-language projector, and 3) a language model.
We select the pretraining dataset that are fully open-source, and have been used in prior works. Specifically, we use the LLaVa v1.5 data mixture, which consists of two subsets used for a multi-stage training pipeline. The first subset consists of a 558K sample mixture of examples sourced from various captioning datasets (e.g., Conceptual Captions, LAION [121]), while the second consists of 665K multimodal instruct tuning examples comprised of synthetic data generated in [122], as well as examples from existing vision-language training sets (e.g., GQA, TextCaps, [123]), and notably, a sample of language-only data from ShareGPT [124].
During the training of our VLM, we adopt a single-stage approach that both the projection and LM are trained while the visual representation modules are frozen. We train for two epochs.
@@ 7.4 CoT Evaluation
7.5 VLM Evaluation
We adopt the evaluation suit in Prismatic VLMs [115] to evaluate the VLM performance. It focuses on evaluations with well-defined metrics, spanning the following three areas:
Open-Ended Visual Question Answering. We evaluate on VizWiz [133] and GQA [123]. VizWiz assess general visual reasoning; VizWiz also contains a series of unanswerable questions. GQA evaluates spatial reasoning.
Localization. Part of the pretraining data mixture contains examples of predicting normalized bounding box coordinates given referring expressions in language. As such, we evaluate bounding box prediction accuracy on RefCOCO, RefCOCO+, and RefCOCOg [134], and on OCID-Ref [135]. RefCOCO focuses on short descriptions with spatial anchors, RefCOCO+ on strictly appearance based descriptions, and RefCOCOg on long, rich descriptions; OCID-Ref is a robotics dataset probing out-of-distribution generalization, with a focus on localizing objects in clutter.
Challenge Sets (Closed-Set Prediction). We evaluate on Visual Spatial Reasoning [119], TallyQA [136], and POPE [137]. VSR consists of challenging True/False questions about individual spatial relationships in diverse scenes; this is an especially challenging task, with most existing models failing to outperform the majority class baseline. TallyQA consists of questions that assess a VLM’s ability to count objects described in language, with expressions that range in complexity. POPE consists of targeted Yes/No questions that assess a VLM’s propensity to hallucinate.
We use the validation sets for all benchmarks except GQA (where use the recommended the test-dev split), VSR (where we use the zero-shot test split), and POPE (where there is only a single evaluation split).
The comparisons with other VLMs are demonstrated in Table 7.5. We compare with LLaVa v1.5 7B. Furthermore, we adopt the same VLM framework and change the LLM backbone to other LLMs such as Llama2 7B and Mistral 7B. Then we train these VLMs and compare with our VLM based on Moxin-7B. We can observe that our model outperforms all other VLM baselines.
[t] Performance comparison for various VLM models. GQA VizWiz RefCOCO+ OCID-Ref VSR POPE TallyQA Ave. LLaVa v1.5 7B (Base) 61.58 54.25 49.47 35.07 51.47 86.57 62.06 57.21 Llama-2 Chat 7B 62.11 56.39 58.5 46.3 61.8 86.8 58.1 61.43 Mistral v0.1 7B 63.3 55.32 65.1 48.8 58.5 87.1 61.7 62.83 Mistral Instruct v0.1 7B 62.71 54.35 64.9 48 57.8 87.5 64.5 62.82 Llama-2 7B 62.44 55.98 59.47 43.89 63.67 86.74 59.22 61.63 Ours 64.88 54.08 71.3 48.4 60.8 87.3 66 64.68
7.6 Moxin Family
@@
[h] Our developed models and their names in our releases. Developed Models Names in Releases Moxin-7B-Enhanced Moxin-7B-LLM Moxin-7B-SFT Moxin-7B-DPO Moxin-7B-Instruct Moxin-7B-DPO-II Moxin-7B-RL-DeepScaleR Moxin-7B-Reasoning Moxin-7B-RL-AReal Moxin-7B-VLM Moxin-7B-VLM
@@ 7.7 Generation Examples
We demonstrate some text generation examples from our Moxin Chat 7B and Llama2 Chat 7B in Table 7.7 and 7.7. Moxin Chat 7B is developed by finetune the Moxin base model on the Tulu2 dataset. The response of Moxin Chat 7B is more accurate and concise.